# SONICS: Synthetic Or Not - Identifying Counterfeit Songs

### v2

@@ 3.2.1 Spectro-Temporal Tokens Transformer
As illustrated in Fig. 3, the proposed SpecTTTra model begins by applying temporal and spectral slicing to copies of the input spectrogram, generating distinct temporal and spectral clips (or patches). These clips are then processed by the Spectro-Temporal Tokenizer, where temporal clips are embedded as temporal tokens $\langle t_{i}\rangle$ and spectral clips as spectral tokens $\langle f_{i}\rangle$, using separate tokenizers. Each token is represented as a vector of shape $\left(1,embed\_dim\right)$. Subsequently, separate positional embeddings are added to the temporal and spectral tokens, as there is no positional relationship between these two types.
@@
It is important to note that while previous work has attempted to utilize both spectral and temporal information, these approaches come with significant limitations. For instance, [31] focuses solely on temporal tokens, neglecting the rich spectral information. On the other hand, methods like [32, 23] employ separate attentions for spectral and temporal information but with ViT-like tokens, which, as discussed in the following section, become highly inefficient and computationally expensive as they dramatically increase with longer audio inputs. In contrast, our approach disentangles spectral and temporal information at the tokenization level and then contextualizes them with attention. This results in a more efficient method compared to existing approaches.
@@ 3.2.2 Spectro-Temporal Tokenization
Traditional Vision Transformers (ViTs) generate grids of patches by simultaneously dividing 2D spectrogram across both temporal and spectral dimensions. If $T$ and $F$ denote the temporal and spectral dimensions of the spectrogram, respectively, and $p$ is the patch size, then the number of patches (or tokens) generated by a ViT ($N_{\nu}$) can be expressed as:
$N_{\nu}=\left(\frac{F}{p}\right)\times\left(\frac{T}{p}\right)=\left(\frac{FT}{p^{2}}\right)$ (1)
This equation highlights the rapid growth in the number of patches as the temporal dimension $T$ increases, leading to a corresponding exponential rise in computational cost. For example, with a spectrogram of size $F=128$ (mel bands) and short audio with $T=128$ (5 sec.), ViT generates $N_{\nu}=64$ tokens (for $p=16$). However, for longer audio with $T=3744$ (120 sec.), the number of tokens surges to $N_{\nu}=1872$, nearly 30 times more than for the shorter audio. Since the computational cost for global attention scales quadratically with the number of tokens, this makes ViTs impractical for long audio classification tasks.
In contrast, the proposed SpecTTTra model adopts an alternative strategy by independently slicing across temporal and spectral dimensions to create separate temporal and spectral patches, as illustrated in Fig. 3. Unlike ViTs, where each patch only has access to limited spectral and temporal information due to simultaneous slicing across both dimensions, SpecTTTra ensures that each temporal patch has access to all spectral components and vice versa. This design leverages the observation that meaningful correlations can exist between distant temporal clips (e.g., between $|t_{0}|$ and $|t_{4}|$, capturing repeated song verses) or between distinct spectral clips (e.g., between $|f_{0}|$ and $|f_{2}|$, differentiating real from fake songs). If $t$ and $f$ represent the sizes of the temporal and spectral clips, respectively, then the number of patches (or tokens) generated by SpecTTTra, denoted as $N_{\psi}$, is given by:
$N_{\psi}$ $=N_{spectral}+N_{temporal}=\left(\frac{F}{f}\right)+\left(\frac{T}{t}\right)$ (2)
Due to the additive nature of this equation, the number of tokens in SpecTTTra grows much more slowly compared to ViTs, thereby keeping computational costs manageable. For instance, using the same parameters as before and setting $t=7$ and $f=5$ in SpecTTTra, the number of tokens for short audio with $T=128$ is $N_{\psi}=43$, and for long audio with $T=3744$, it increases to $N_{\psi}=560$, which is approximately 3.4 times fewer than in the ViT model. This substantial reduction in token count makes SpecTTTra significantly more computationally efficient. We further explore three variants of SpecTTTra, differentiated by the sizes of their temporal ($t$) and spectral ($f$) patches: SpecTTTra-$\alpha$ ($f=1$, $t=3$), SpecTTTra-$\beta$ ($f=3$, $t=5$), and SpecTTTra-$\gamma$ ($f=5$, $t=7$). Fig. 4 illustrates the rate at which the number of tokens increases for different SpecTTTra variants and ViTs as the audio length (number of time frames) grows.
@@
The resulting spectral and temporal clips are processed by the Spectro-Temporal Tokenizer (STT) to create spectral and temporal tokens, respectively. The proposed STT block consists of separate Spectral and Temporal Tokenizers, which share identical architecture but differ in their objectives based on their input. These tokenizers use a Linear layer to map the spectral/temporal clips to spectral/temporal tokens, followed by GELU activation, addition of learnable positional embeddings similar to ViT, and finally, layer normalization. For efficiency, the slicing and tokenization operations are merged using 1D convolution. Given an input spectrogram $x\in\mathbb{R}^{F\times T}$, where $T$ and $F$ denote the temporal and spectral dimensions of the input spectrogram, the mathematical expression for tokenization is as follows:
@@
$\langle t\rangle$ $=\text{LayerNorm}\left(\text{GELU}\left(x^{\top}_{t}\right)+\hat{p}_{t}\right)$
@@
$\langle f\rangle$ $=\text{LayerNorm}\left(\text{GELU}\left(x^{\top}_{f}\right)+\hat{p}_{f}\right)$
@@
Here, $\langle t\rangle\in\mathbb{R}^{\frac{T}{t}\times D}$ and $\langle f\rangle\in\mathbb{R}^{\frac{F}{f}\times D}$ denote the temporal and spectral tokens, where $t$ and $f$ are the temporal and spectral clip sizes, respectively, and $D$ is the token embedding dimension. The parameters $c_{i}$, $c_{o}$, $k$, and $s$ represent the input channels, output channels, kernel size, and stride size of the CNN layers. Additionally, $\hat{p}$ and $\top$ denote the learnable positional embeddings and Transpose operation. Further implementation details can be found in the Appendix.
@@
5 ConvNeXt 0.99 0.99 0.62 0.62 0.99 0.99 0.99 0.90 0.82 0.80 0.90 0.82 0.98
ViT 0.98 0.95 0.79 0.57 0.86 0.79 0.79 0.92 0.78 0.76 0.79 0.80 0.79
SpecTTTra-$\gamma$ 0.99 0.98 0.51 0.10 0.99 0.98 0.97 0.87 0.61 0.62 0.76 0.63 0.98
SpecTTTra-$\beta$ 0.99 0.98 0.61 0.18 0.99 0.95 0.94 0.89 0.66 0.66 0.78 0.69 0.94
SpecTTTra-$\alpha$ 0.99 0.99 0.68 0.26 0.99 0.93 0.92 0.91 0.69 0.70 0.80 0.71 0.92
120 ConvNeXt 0.99 0.99 0.77 0.95 1.0 0.98 0.98 0.94 0.95 0.93
0.96
0.95 0.98
ViT 1.00 0.99 0.82 0.53 0.99 0.99 0.98 0.95 0.80 0.80 0.89 0.82 0.98
SpecTTTra-$\gamma$ 0.99 0.99 0.83 0.59 0.99 0.99 0.98 0.95 0.82 0.81 0.91 0.83 0.98
SpecTTTra-$\beta$ 1.00 0.99 0.87 0.62 0.99 0.99 0.99 0.96 0.84 0.82 0.92 0.86 0.99
SpecTTTra-$\alpha$ 1.00 0.99 0.88 0.71 0.99 0.99 0.99 0.98 0.89 0.87 0.94 0.91 0.99
Table 3: AI Benchmark
4 Experiments and Results
4.1 Dataset
We conduct all experiments using the proposed SONICS dataset, which is divided into train, valid, and test sets. To ensure comprehensive evaluation, the valid and test sets include cases with unseen algorithms (e.g., Suno v2, Suno v3, Udio 32) and new singers. We also prevent data leakage by ensuring that song pairs from the same (lyrics, style) inputs are exclusively in either the training or valid-test sets, not in both. The distribution of the train and test sets is shown in Table 2, with the valid set distribution available in the Appendix.
4.2 Implementation Details
We conduct model training on an NVIDIA A6000 GPU with 48GB RAM, using WandB [3] for tracking. Generating the SONICS dataset costs \$1,055, allocated across GPT-4o (\$405), Gemini 1.5 Flash (\$80), Suno (\$390), and Udio (\$180). We generate spectrograms with a 16,000 Hz sample rate, n_fft = win_length = 2048, hop_length = 512, and n_mels = 128, yielding a $128\times 128$ spectrogram for 5 sec and $128\times 3744$ for 120 sec audio. Any song shorter than input length is zero-padded randomly, while for longer songs, a random crop is used. However, during the test, to maintain determinism, padding is done on the right side and cropped segments are taken from the middle. We use small and tiny variants of ViT (patch size = 16) and ConvNeXt from the timm [29] library. In SpecTTTra, we use the same model configuration as ViT-small. It is noteworthy that we trained all models from scratch for fair comparison. We train for 50 epochs using Binary Cross-Entropy loss with 0.02 label smoothing and a batch size of 128. Optimization is performed with AdamW and a cosine learning rate scheduler from timm, including a 5-epoch warm-up and an initial learning rate of 0.0005. While existing methods [34, 33, 30] typically use Equal Error Rate (EER) as a metric, we prioritize the F1 score (binary average, threshold = 0.5) due to EER’s susceptibility to class imbalance. We also evaluate Sensitivity (Sens.) and Specificity (Spec.) to assess performance across fake and real classes. Further details can be found in the Appendix.
4.3 Benchmarks
4.3.1 AI Benchmark
The comparative analysis of the proposed SpecTTTra models against ViT and ConvNeXt models is presented in Table 3. The results reveal a significant performance gain (6% for ConvNeXt, 10% for ViT, and 14% for SpecTTTra-$\alpha$) in the Overall F1 score when using long songs. This finding substantiates our claim that leveraging long-context information is crucial for enhancing fake song detection. Furthermore, the advantage of longer audio duration is more prevalent in transformer-based models (ViT or SpecTTTra variants) compared to CNN models (ConvNeXt). This can be attributed to the global attention mechanism in transformer models, which effectively captures long-range dependencies within the input data. However, despite the absence of global attention, ConvNeXt outperforms both ViT and SpecTTTra models in both short and long audio scenarios. We hypothesize that this is due to the inherent inductive biases present in CNNs, which are lacking in transformers, leading the latter to require larger datasets to achieve comparable performance [17, 7]. Notably, the proposed SpecTTTra-$\alpha$, while trailing ConvNeXt by 10% in the F1 score for short audio, narrows this gap to 4% for long audio, highlighting its potential in long-context detection tasks. Another intriguing observation is the performance of ViT, which, despite its large number of tokens (or patches), is outperformed by all SpecTTTra variants in terms of overall F1 score for long audio, solidifying SpecTTTra’s effectiveness. This could be due to an overload of redundant information from ViT’s numerous patches, which may not contribute effectively to the detection task. Moreover, an interesting pattern can be observed across all models – real songs are more easily identified than fake ones, as indicated by higher specificity and lower sensitivity scores.
Diving deeper into different partitions of test data, we observe that all detection models achieve better performance on seen algorithms (Suno v3.5 and Udio 130) compared to unseen ones (Suno v2, Suno v3, and Udio 32). Particularly, they struggle more with the Udio algorithms, with the most pronounced difficulty observed for Udio 32. However, ConvNeXt performs relatively well in detecting the Udio 32 algorithm, achieving a sensitivity of 95%, possibly due to the CNN’s strong locality. Interestingly, despite being an unseen algorithm, the detectors perform comparably well on Suno v3 as they do on the seen Suno v3.5 algorithm, suggesting a possible algorithmic similarity between the two. On the other hand, for short audio samples, the models perform better on songs with seen speakers than those with unseen speakers, a gap that diminishes when longer audio is used. Finally, in fake type partitions, all detectors excel in detecting HF songs, which can be linked to the exclusive presence of Suno algorithms, where the detectors generally perform better compared to Udio algorithms. Among MF and FF songs, the models show slightly lower performance on FF songs, which might be due to the diverse unseen combination of topics, moods, and genres that differ more significantly than other fake songs.

### v4

@@ 3.2.1 Spectro-Temporal Tokens Transformer (SpecTTTra)
As illustrated in Fig. 3, the proposed SpecTTTra model begins by applying temporal and spectral slicing to copies of the input spectrogram, generating distinct temporal and spectral clips/patches indicated by $|t_{i}|$ and $|f_{i}|$ respectively where $i$ denotes the position of each clip. These clips are then processed by the Spectro-Temporal Tokenizer, where temporal clips are embedded as temporal tokens $\langle t_{i}\rangle$ and spectral clips as spectral tokens $\langle f_{i}\rangle$, using separate tokenizers, where each token is represented as a vector of shape $\left(1,embed\_dim\right)$. Subsequently, separate positional embeddings ($p_{i}^{t}$ for temporal and $p_{i}^{f}$ for spectral) are added to the respective tokens. Separate embeddings are used because there is no positional relationship between temporal and spectral tokens.
@@
It is important to note that while previous work has attempted to utilize both spectral and temporal information, these approaches come with significant limitations. For instance, Yadav et al. (2023); Gong et al. (2021) focuses solely on temporal tokens, neglecting the rich spectral information. On the other hand, methods like Zadeh et al. (2019); Shul & Choi (2024) employ separate attentions for spectral and temporal information but with ViT-like tokens, which, as discussed in the following section, become highly inefficient and computationally expensive as they rapidly increase with longer audio inputs. In contrast, our approach disentangles spectral and temporal information at the tokenization level and later contextualizes them with attention, leading to greater efficiency.
@@ 3.2.2 Spectro-Temporal Tokenization
The ViT model creates grids of small square patches by simultaneously dividing a 2D spectrogram along both temporal and spectral dimensions, resulting in each patch having access to only limited spectral and temporal information. More critically, as indicated in Eq. 1, the number of patches/tokens ($N_{\nu}$) increases rapidly with the temporal dimension $T$. For example, for a spectrogram of size $F=128$ (spectral dimension) with a short audio duration of 5 sec ($T=128$) , ViT (patch size, $p=16$) generates tokens $N_{\nu}=64$. However, for longer audio with 120 sec ($T=3744$), the number of tokens surges to $N_{\nu}=1872$, nearly 30x more than for the shorter audio. Since the computational cost for global attention in ViT scales quadratically with the number of tokens, this makes it less practical for long audio classification. Similarly, EfficientViT, despite having linear global attention, multi-scale operations coupled with numerous tokens still result in high complexity.
Equation for ViT Token Count
If $T$ and $F$ denote the temporal and spectral dimensions of the spectrogram, respectively, and $p$ is the patch size, then the number of patches (or tokens) generated by a ViT ($N_{\nu}$) can be expressed as: $N_{\nu}$ $=\left(\frac{F}{p}\right)\times\left(\frac{T}{p}\right)$ (1)
Equation for SpecTTTra Token Count
If $t$ and $f$ represent the sizes of the temporal and spectral clips, respectively, then the number of patches (or tokens) generated by SpecTTTra, denoted as $N_{\psi}$, is given by: $N_{\psi}$ $=N_{spectral}+N_{temporal}$ (2) $=\left(\frac{F}{f}\right)+\left(\frac{T}{t}\right)$ (2)
@@
In contrast, SpecTTTra performs slicing independently for temporal and spectral dimensions, as shown in Fig. 3. This approach ensures that each temporal patch has access to the full spectral information, and each spectral patch contains the complete temporal information. This design leverages the observation that meaningful correlations can exist between distant temporal clips (e.g., between $|t_{0}|$ and $|t_{4}|$, capturing repeated song verses) or between distinct spectral clips (e.g., between $|f_{0}|$ and $|f_{2}|$, capturing harmonics). Importantly, as demonstrated in Eq. 2, due to the additive nature of token generation in SpecTTTra, the number of tokens grows more slowly compared to ViTs, significantly reducing computational costs. For instance, using the same parameters as before and setting the temporal and spectral clip sizes to $t=7$ and $f=5$, the number of tokens in SpecTTTra for short audio is $N_{\psi}=43$, and for long audio, it increases to $N_{\psi}=560$, which is approximately 3.4x fewer than in the ViT model. This substantial reduction in token count makes SpecTTTra significantly more computationally efficient. We further explore three variants of SpecTTTra, differentiated by the sizes of their temporal ($t$) and spectral ($f$) patches: SpecTTTra-$\alpha$ ($f=1$, $t=3$), SpecTTTra-$\beta$ ($f=3$, $t=5$), and SpecTTTra-$\gamma$ ($f=5$, $t=7$). Fig. 4 illustrates the rate at which the number of tokens increases for different SpecTTTra variants and ViT as the audio length grows.
The resulting spectral and temporal clips are processed by the Spectro-Temporal Tokenizer (STT) to create spectral and temporal tokens, respectively. The proposed STT block consists of separate Spectral and Temporal Tokenizers, which share identical architecture but differ in their objectives based on their input. These tokenizers use a Linear layer to map the spectral/temporal clips to spectral/temporal tokens, followed by GELU activation, addition of learnable positional embeddings similar to ViT, and finally, layer normalization. For efficiency, during implementation, the slicing and tokenization operations are merged using 1D CNN layer. Given an input spectrogram $x\in\mathbb{R}^{F\times T}$, the mathematical expression for tokenization is as follows:
@@
$\langle t\rangle$ $=\text{LayerNorm}\left(\text{GELU}\left(x^{\top}_{t}\right)+\hat{p}^{t}\right)$
@@
$\langle f\rangle$ $=\text{LayerNorm}\left(\text{GELU}\left(x^{\top}_{f}\right)+\hat{p}^{f}\right)$
@@
Here, $\langle t\rangle\in\mathbb{R}^{\frac{T}{t}\times D}$ and $\langle f\rangle\in\mathbb{R}^{\frac{F}{f}\times D}$ denote the temporal and spectral tokens, where $D$ is the token embedding dimension. The parameters $c_{i}$, $c_{o}$, $k$, and $s$ represent the input channels, output channels, kernel size, and stride size of the CNN layers. Additionally, $\hat{p}$ and $\top$ denote the learnable positional embeddings and Transpose operation. Further implementation details can be found in the Appendix.
4 Experiments and Results
4.1 Dataset
We conduct all experiments using the proposed SONICS dataset, which is divided into train, valid, and test sets. To ensure comprehensive evaluation, the valid and test sets include cases with unseen algorithms (e.g., Suno v2, Suno v3, Udio 32) and unseen singers. We also prevent data leakage by ensuring that song pairs from the same (lyrics, style) inputs are exclusively in either the training or valid-test sets, not in both. The distribution of the train, test, and valid sets is shown in Table 3.
4.2 Implementation Details
To train models, we resampled both real and fake songs to 16kHz and generated spectrograms with n_fft = win_length = 2048, hop_length = 512, and n_mels = 128, yielding a $128\times 128$ spectrogram for 5 sec and $128\times 3744$ for 120 sec audio. Any song shorter than input length is zero-padded randomly, while for longer songs, a random crop is used. We also apply MixUp (Zhang, 2017) and SpecAugment (Park et al., 2019) augmentations during training to improve generalization. However, during the test, to maintain determinism, padding is done on the right side and cropped segments are taken from the middle. We conduct our training on an NVIDIA A6000 GPU with 48GB RAM, using WandB for tracking. We use ViT-small (patch size = 16) and ConvNeXt-tiny along with EfficientViT-B2 from the timm (Wightman, 2019) library. In SpecTTTra, we use the same model configuration as ViT-small. We train all models for 50 epochs from scratch using Binary Cross-Entropy loss with 0.02 label smoothing (Szegedy et al., 2016). Optimization is performed with AdamW (Loshchilov, 2017) and a cosine learning rate scheduler from timm, including a 5-epoch warm-up. While existing methods (Zang et al., 2024b; a; Xie et al., 2024) use Equal Error Rate (EER) as a metric, we prioritize the F1 score (binary average, threshold = 0.5) as our primary metric due to EER’s susceptibility to class imbalance. We also evaluate Sensitivity (Sens.) and Specificity (Spec.) to assess performance across fake and real classes.
4.3 Benchmarks
4.3.1 AI Benchmark
The comparative analysis of the proposed SpecTTTra models against other existing models is presented in Table 4. The results reveal a significant performance gain (6% for ConvNeXt, 8% for EfficientViT, 10% for ViT, and 17% for SpecTTTra-$\alpha$) in the overall F1 score when using long songs. This finding substantiates our claim that leveraging long-context information is crucial for enhancing fake song detection. Additionally, the advantage of longer audio duration is more prevalent in transformer-based models such as ViT and SpecTTTra, as well as in the hybrid EfficientViT model, compared to the CNN-based ConvNeXt. Notably, the proposed SpecTTTra-$\alpha$, while trailing ConvNeXt by 10% in the F1 score for short audio, outperforms it in long audio. This can be attributed to the global attention mechanism in transformer models, which effectively captures long-range dependencies within the input data. In contrast, models with CNN components tend to perform better on shorter audio. Specifically, ConvNeXt and EfficientViT achieve overall F1 scores of 90% and 87%, respectively, outperforming all transformer-based models on short audio. However, despite the absence of global attention, ConvNeXt demonstrates competitive performance compared to SpecTTTra-$\alpha$ on long audio and outperforms ViT, EfficientViT, and other SpecTTTra variants in both short and long audio scenarios. We hypothesize that this is due to the inherent inductive biases present in CNNs, which are lacking in transformers, leading the latter to require larger datasets to reach their true potential (Liu et al., 2021; Dosovitskiy et al., 2020). Another intriguing observation is the performance of ViT, which, despite its large number of tokens (or patches), is outperformed by the $\alpha$ and $\beta$ variants of SpecTTTra and is only on par with the $\gamma$ variant in terms of overall F1 score for long audio, reinforcing SpecTTTra’s effectiveness. We hypothesize that this is due to an overload of redundant information from ViT’s numerous patches, which may not contribute effectively to the detection task. Moreover, it can also be observed across all models that real songs are more easily identified than fake ones, as indicated by higher specificity and lower sensitivity scores.
Diving deeper into different partitions of test data, we observe that all detection models achieve better performance on seen algorithms (Suno v3.5 and Udio 130) compared to unseen ones (Suno v2, Suno v3, and Udio 32). Particularly, they struggle more with the Udio algorithms, with the most pronounced difficulty observed for Udio 32. However, ConvNeXt and SpecTTTra-$\alpha$ perform relatively well in detecting the Udio 32 algorithm, achieving a sensitivity of 96% and 95% respectively. Interestingly, despite being an unseen algorithm, the detectors perform comparably well on Suno v3 as they do on the seen Suno v3.5 algorithm, suggesting a possible algorithmic similarity between the two. Conversely, for short audio samples, the detectors perform slightly better on songs with seen speakers than those with unseen speakers, a gap that diminishes when longer audio is used. Finally, in Fake Type partitions, all detectors excel in identifying HF songs, likely due to the exclusive presence of Suno algorithms, where detectors generally perform better compared to Udio algorithms. Among MF and FF songs, the models exhibit a slightly lower performance pattern on FF songs.
Table 4: Performance comparison of SpecTTTra and conventional AI models on varying audio lengths, with F1 score as the primary evaluation metric. Here Real/Human and Fake/AI songs denoting Negative and Positive classes, respectively. $\dagger$ indicates unseen algorithms during training.
@@ 4.3.1 AI Benchmark
5 ConvNeXt 0.62 0.99 0.99 0.62 0.99 0.99 0.99 0.90 0.82 0.80 0.90 0.82 0.98
ViT 0.79 0.95 0.98 0.57 0.86 0.79 0.79 0.92 0.78 0.76 0.79 0.80 0.79
EfficientViT 0.66 0.98 0.99 0.49 0.97 0.99 0.98 0.90 0.76 0.74 0.87 0.78 0.98
SpecTTTra-$\gamma$ 0.51 0.98 0.99 0.10 0.99 0.98 0.97 0.87 0.61 0.62 0.76 0.63 0.98
SpecTTTra-$\beta$ 0.61 0.98 0.99 0.18 0.99 0.95 0.94 0.89 0.66 0.66 0.78 0.69 0.94
SpecTTTra-$\alpha$ 0.68 0.99 0.99 0.26 0.99 0.93 0.92 0.91 0.69 0.70 0.80 0.71 0.92
120 ConvNeXt 0.77 0.99 0.99 0.95 1.00 0.98 0.98 0.94 0.95 0.93 0.96 0.95 0.98
ViT 0.82 0.99 1.00 0.53 0.99 0.99 0.98 0.95 0.80 0.80 0.89 0.82 0.98
EfficientViT 0.73 0.98 1.00 0.95 1.00 0.97 0.97 0.92 0.92 0.94 0.95 0.94 0.97
SpecTTTra-$\gamma$ 0.98 0.99 1.00 0.37 1.00 0.99 0.99 0.99 0.77 0.76 0.88 0.79 0.99
SpecTTTra-$\beta$ 0.87 0.99 1.00 0.62 0.99 0.99 0.99 0.96 0.84 0.82 0.92 0.86 0.99
SpecTTTra-$\alpha$ 0.78 0.99 1.00 0.96 1.00 0.99 0.99 0.98 0.89 0.87 0.97 0.96 0.99